OpenStudy (anonymous):

factor completely: 16x^4-16

mathslover (mathslover):

Hint: \[\large{a^2-b^2=(a+b)(a-b)}\] \[\large{(16x^4)-(16)}\] \[\large{\implies (4x^2)^2-(4)^2}\] hence put a = \(4x^2\) and b = \(4\)

mathslover (mathslover):

@Mberdeja501 let me know that what you get after following my instructions.....

OpenStudy (anonymous):

(4x^2+2)(4x^2-2)?

mathslover (mathslover):

b = 4 ... correct it again... You had put 2 there instead of 4

OpenStudy (anonymous):

(4x^2+4)(4x^2-4)?

mathslover (mathslover):

good.. Now, see the second bracket : \(\large{(4x^2-4) = [{(2x)^2-(2)^2}}]\) can you again factorize this into (a+b)(a-b) ?

mathslover (mathslover):

@Mberdeja501 let me know what you get : after following the above expressions

mathslover (mathslover):

@Mberdeja501 ... r u stuck anywhere ? let me know where r u having problem?

OpenStudy (anonymous):

so then i get 2(2x^2+2)(x-1)(x+1)

mathslover (mathslover):

\[\large{(4x^2+4)(4x^2-4) \implies 4(x^2+1)\times 4(x^2-1)}\] \[\large{16(x^2+1)(x^2-1) \space \textbf{Factorize x^2-1 now} }\] \[\large{16(x^2+1)(x+1)(x-1) \space \textbf{(Using a^2-b^2=(a+b)(a-b) identity)}}\]

mathslover (mathslover):

got it?

mathslover (mathslover):

@Mberdeja501 ?

OpenStudy (anonymous):

uhhhhhmmm, totally lost me

mathslover (mathslover):

we had this : \[\large{(4x^2+4) \times (4x^2-4)}\] correct?